Shift summary for owl 10/7/99

Philippe Gouffon (pgouffon@fdei01)
Thu, 07 Oct 1999 07:02:40 -0500

Shift Summary for Owl Shift, 10/7/99

On Shift: Philippe Gouffon

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

========================================================================
Overview of shift:
No beam. TeV purification in progress. Nothing going on at KTeV.

========================================================================
Description of Shift:

02:50 Doug stops by to get his battery charger

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Doug recharged the battery for the EOM used to measure Z'. When an
appropriate number of people are available, re-install and check z position

------- End of Forwarded Message